OpenStudy (anonymous):

find the average rate of change of the function from x1, to x2 f(x)= √2x from x1=2 to x2=8 Help pleaseeeee

OpenStudy (dumbcow):

its basically slope formula \[\frac{f(8) -f(2)}{8-2}\]

OpenStudy (anonymous):

so what I don't get is how did they get 1/3 as the answer???

OpenStudy (dumbcow):

i dont get it either :) answer should be 1/2 \[f(x) = \sqrt{2x}\] right

OpenStudy (dumbcow):

oh oops, yeah its 1/3

OpenStudy (dumbcow):

f(8) = sqrt16 = 4 f(2) = sqrt4 = 2 \[\frac{4-2}{8-2} = \frac{2}{6}=\frac{1}{3}\]

OpenStudy (anonymous):

sorry wait from where did you get the 4

OpenStudy (dumbcow):

f(8) = sqrt(2*8) = sqrt16 = 4